The initial partial pressures are PA = 0.000 torr, PB = 0.000 torr, and

PC = 150. torr, and the total pressure at equilibrium is 200.0 torr. What are

the values for Kp and Kc at 800. K temperature?


1
Expert's answer
2021-03-25T04:08:19-0400

A(g) + B(g) "\leftrightarrow" C(g)

If total pressure is 200 torr, then PA = x torr, PB = x torr, PC = (150 - x) torr

200 = 2x + (150 - x)

x = 50

Kp = "\dfrac{P_C}{P_AP_B}" ="\dfrac{100}{50\times50}" = 0.04 (torr-1)

Kc = "\dfrac{Kp}{(RT)^{\Delta n}}" ="\dfrac{0.04}{(8.134\times800)^{-1}}" = 266.05